Solve the following quadratic equations by factoring or by any of the techniques
x^2 - 9x - 4 =6

Thanks

Answer by checkley71(8403)

X~2-9X-4=6 OR X~2-9X-10=0 OR (X-10)(X+1)=0 OR X-10=0 OR X=10 & X+1=0 OR X=-1